OpenStudy (anonymous):

Show that TanO*SinO/1-CosO=1+1/CosO

OpenStudy (anonymous):

you can replace O with x, as it can be thought as Zero (0)..

OpenStudy (anonymous):

\[\frac{ \tan \Theta \sin \Theta }{ 1-\cos \Theta? }= 1+\frac{ 1 }{ \cos \Theta }\]

OpenStudy (anonymous):

\[\frac{\sin(\theta) \cdot \tan(\theta)}{1 - \cos(\theta)}\]

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

Start by : \[\tan(\theta) =\frac{\sin(\theta)}{\cos(\theta)}\]

OpenStudy (anonymous):

What do you get when you will use this??

OpenStudy (anonymous):

\[\frac{ \sin ^{2\Theta} }{ 1-\cos ^{2} \Theta }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

LHS: \[\implies \frac{\sin^2(\theta)}{\cos(\theta) [1 - \cos(\theta)]}\]

OpenStudy (anonymous):

No, the cos(theta) will be multiplied by 1 also, you forgot to do that.. Don't multiply cos(theta) inside the brackets..

OpenStudy (anonymous):

\[1 + \cos(\theta) = 2 \cos^2(\frac{\theta}{2})\\1 - \cos(\theta) = 2\sin^2(\frac{\theta}{2})\]

OpenStudy (anonymous):

For the 1-cos(theta) in denominator, use the above identity..

OpenStudy (anonymous):

\[\frac{\sin^2(\theta)}{\cos(\theta)[2 \sin^2(\frac{\theta}{2})]}\]

OpenStudy (anonymous):

Now you also know that: \[\sin(\theta) = 2 \sin(\frac{\theta}{2})\cos(\frac{\theta}{2})\]

OpenStudy (mathmath333):

u can also use\[\sin^2 \theta(1+\cos \theta)/\cos \theta(1-\cos \theta)(1+\cos \theta)\]

OpenStudy (anonymous):

\[\frac{4 \sin^2(\frac{\theta}{2}) \cos^2(\frac{\theta}{2})}{2 \sin^2(\frac{\theta}{2}) \cos(\theta)} \\ \\\]

OpenStudy (anonymous):

\[\implies \frac{2\cos^2(\frac{\theta}{2})}{\cos(\theta)} \implies \frac{1 + \cos(\theta)}{\cos(\theta)} \implies \color{blue}{ 1 + \frac{1}{\cos(\theta)}}\]

OpenStudy (anonymous):

Oh, I did not see that, you can also try rationalising the denominator..

OpenStudy (anonymous):

@mathmath333 method is simpler one..

OpenStudy (anonymous):

\[LHS \implies \frac{\sin^2(x)}{\cos(x)[1 - \cos(x)]}\] I am using x for \(\theta\) here,,

OpenStudy (anonymous):

Here multiply and divide by (1 + cos(x)).. \[\frac{\sin^2(x) [1 + \cos(x)]}{\cos(x) (1 - \cos(x))[1 + \cos(x)]} \implies \frac{\sin^2(x) [1 + \cos(x)]}{\cos(x) \cdot \sin^2(x)} \\ \frac{1 + \cos(x)}{\cos(x)} \implies 1 + \frac{1}{\cos(x)}\]

OpenStudy (anonymous):

Note that: In denominator I have done: \[(1 - \cos(x))(1 + \cos(x)) = 1 - \cos^2(x) = \sin^2(x)\]

OpenStudy (anonymous):

Based on what did we multiply and divide by (1+cos(x))?

OpenStudy (anonymous):

Logical thinking.. As you have (1- cos(x)) on down, so your mind must tell you that if I multiply it with (1+ cos(x)), you will get sin^2(x) in denominator, and upper sin(x) can get cancelled with it,..

OpenStudy (anonymous):

In proving Trigonometric Identities, just start from the identity you know, there is nothing hard and fast rule, from all the methods you will get the same answer, some method can take longer steps like MINE, and some can take smaller like @mathmath333

OpenStudy (anonymous):

Just start by what you know, you will definitely reach to your answer, but while proving, do look at your RHS, as you want to make your LHS equals RHS.. So, you can't do anything that will take away from your RHS.. You will get it, solve like 50 of these problems, you will gain confidence.. :)

OpenStudy (anonymous):

I gotta go now.. @mathmath333 will take on now onwards.. :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (mathmath333):

any problem @Attyeh

OpenStudy (mathmath333):

for expertise in trignometry u should remember and try to prove all trignometric identities ,formulas

OpenStudy (anonymous):

I do face a problem regarding trigonometry, do you have a list of the most needed identities? Im doin P1 in October and im trying to Ace it so any help would be appreciated @mathmath333